Assume the specifications for a television remote control part (in inches) are 10.00 ± 0.10, and the Taguchi loss function is estimated to be L(x) = 15,500 (x - T)2. Determine the estimated loss per part if the quality characteristic (i.e., actual process performance) under study takes on a value of 9.90 inches.

Question 6 options:

a). More than $50 but less than or equal to $100

b). More than $100 but less than or equal to $150

c). More than $150 but less than or equal to $200

d). More than $200
Answer #1

L(x) = k(x - T)2 = 15500( 9.9 - 10)2 = $155

SO, c is the answer.

Note: I tried my best in answering the question. It'd be appreciated if you leave an upvote. Thanks in advance.